# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ XI Future Research Areas
  • Advanced AE architectures: investigate more advanced AE architectures, such as bayesian AEs [172], diffusion AEs [173], contractive AEs [174], denoising AEs [175], semi-supervised AEs [176], and masked AEs [177]. Combining and modifying these architectures could lead to improved performance and new applications in communication systems.
@@
  • Hybrid models: it develops hybrid models that combine the strengths of both data-driven and model-based approaches. This could involve integrating traditional mathematical models with deep learning techniques to enhance the robustness and adaptability of communication systems.
  • Quantitative comparisons: it conducts quantitative comparisons of system performance between different AE architectures. This includes evaluating complexity, efficiency, and effectiveness in various communication scenarios, as well as comparing hardware implementations such as FPGA.
  • Handling high computational complexity: focusses on optimising communication systems for resource-constrained devices, such as FPGAs. This includes developing low-complexity AE architectures and exploring techniques such as low-precision compression and binary neural networks.
@@
  • Non-differentiable channels: it explores alternative approaches for end-to-end learning in real-world scenarios where the channel and transceiver are non-differentiable or the channel is unknown. This could involve reinforcement learning, stochastic perturbation techniques, gradient-free training methods, and genetic algorithms.
  • DSP techniques: it designs AEs for specific DSP techniques, such as equalisation, CPE, and LPN compensation. This includes integrating differentiable DSP algorithms into AE models for end-to-end optimisation.
  • Physical limitation: future research should focus more on integrating physical limitations into end-to-end learning frameworks. Although minimising the reconstruction loss is crucial, solely focussing on this objective may not always yield optimal latent representations. Incorporating physical constraints such as channel capacity, power budget, and bandwidth limitations is essential to ensure practical feasibility. However, excessive imposition of constraints can potentially hinder reconstruction performance due to noise in the received data. Therefore, carefully balancing the trade-off between reconstruction loss and physical constraints is vital to achieve optimal system optimisation.
  • Semantic communication: it enhances AE-based semantic communication systems to improve the reliability of conveying meaning over traditional data recovery methods. This includes developing new metrics for evaluating semantic noise and exploring conceptual spaces for modelling meaning geometrically.
  • Quantum communication: it investigates the combination of quantum and classical approaches to improve end-to-end communication systems. This includes leveraging QNNs to improve robustness against signal degradation and exploring hybrid quantum-classical AE architectures. Furthermore, adapting AEs for purely quantum communication systems is a critical research area.
  • Real-world deployment: it addresses the challenges of real-world deployment, such as the need for extensive training data, the risk of overfitting, and the complexity of practical implementations. This includes developing robust training strategies, data augmentation techniques, and methods for handling noisy or corrupted data.
XII Conclusion
@@
The integration of autoencoders (AEs) into communication systems represents a paradigm shift from traditional model-based approaches to data-driven methodologies. By addressing the limitations of traditional mathematical models, this survey highlighted the potential of AEs to overcome the inherent limitations of mathematical models, providing more accurate and adaptable solutions for next-generation communication systems. The review also highlighted the extensive application of AEs in various domains, including wireless, optical, semantic, and quantum communications, and categorised their contributions in areas such as transceiver design, channel modelling, and digital signal processing. Despite the progress, challenges persist, such as handling non-differentiable channels, computational complexity, and overfitting. Despite promising results, several challenges remain, including the need for extensive training data, the risk of overfitting, and the complexity of real-world deployment. Moreover, this paper conducted a thorough examination of the computational complexity associated with AE-based systems using floating-point operations per second (FLOPS), facilitating a quantitative evaluation of their efficiency and scalability. This insight is crucial to guide the practical implementation of AEs in resource-constrained environments, paving the way for their adoption in real-world communication systems. Future research should focus on addressing these challenges, exploring advanced AE architectures, and developing hybrid models that combine the strengths of both data-driven and model-based approaches. In doing so, the field can move closer to realising the full potential of AEs in creating efficient, reliable, and high-performance communication systems.
Appendix A Counting FLOPs for Activation Functions
A-A ReLU
The calculation of FLOPs for the ReLU activation function is based on the fact that ReLU performs a simple comparison operation for each neurone. Specifically, ReLU computes
$\text{ReLU}(x)=\max(0,x),$ (A.1)
which involves one comparison per neuron. Therefore, for a layer with $N$ neurones, the number of FLOPs for the ReLU activation function is $N$ FLOPs.
A-B ELU
The ELU is more complex than the ReLU. It involves several operations for each neurone. The ELU function is defined as,
$\text{ELU}(x)=\begin{cases}x&\text{if }x\geq 0\\ \alpha(e^{x}-1)&\text{if }x<0\end{cases}.$ (A.2)
The number of FLOPs for the ELU function is calculated as: 1 FLOP for comparison $x\geq 0$, usually around 5 FLOPs for exponential $e^{x}$, 1 FLOP for subtraction, and 1 FLOP for multiplication. Thus, it requires approximately 8 FLOPs for each neurone. This leads to a total of $8N$ for a layer with $N$ neurones.
A-C Leaky ReLU
It is a variant of the ReLU function designed to address the problem of dying ReLU by allowing a small non-zero gradient $\alpha$ when the input is negative. The Leaky ReLU function is defined as:
Leaky $\text{Leaky ReLU}(x)=\begin{cases}x&\text{if }x\geq 0\\ \alpha x&\text{if }x<0\end{cases}.$ (A.3)
For each neurone, 1 FLOP for comparison to check if $x\geq 0$ and 1 FLOP for the multiplication of $\alpha x$ when $x<0$. Thus, the total number of FLOPs for the Leaky ReLU activation function is $2N$ FLOPs for a layer with $N$ neurones.
A-D Tanh
The Tanh activation function is defined as:
$\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}.$ (A.4)
Typically around 5 FLOPs are required for one Exponential Operation, 1 FLOP for subtraction, 1 FLOP for addition, and 1 FLOP for division. This results in 11 FLOPs for the Tanh activation function per neurone. For a layer with $N$ neurones, it is required to have $13N$ FLOPs.
A-E SELU
The SELU activation function is defined as:
$f(x)=\begin{cases}\lambda x&\text{if }x>0\\ \lambda\alpha(e^{x}-1)&\text{if }x\leq 0\end{cases},$ (A.5)
where $\lambda$ and $\alpha$ are predefined constants. For each neurone, one needs 1 FLOP for comparison to check if $x\geq 0$, 5 FLOPs for exponential operation, 1 FLOP for subtraction, and 2 FLOPs for the two multiplication operations. The SELU activation function involves $9N$ FLOPs for a layer with $N$ neurones.
A-F Swish
The Swish activation function is defined as:
$\text{Swish}(x)=x\cdot\sigma(x)$ (A.6)
where $\sigma(x)$ is the sigmoid function as $\sigma(x)=\frac{1}{1+e^{-x}}$.
Sigmoid function involves 5 FLOPs for exponential operation, 1 FLOP for addition and 1 FLOP for subtraction. The Swish requires 1 FLOP for multiplication. This results in $8N$ FLOPs for the Swish activation function per layer with $N$ neurones.
A-G Softmax
The Softmax activation function is commonly used in the output layer of neural networks. It converts raw output (logits) into probabilities, ensuring that the sum of the probabilities for each class is 1. It is defined as:
$\text{Softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$ (A.7)
where $x\_i$ is the input to the $i$-th neuron, and the denominator is the sum of exponentials of all inputs. It involves 5 FLOPs for exponential per input, $N-1$ FLOPs for summation, and 1 FLOP for division per input. Therefore, for an output layer with $N$ neurones, the total FLOP is $15N+(N-1)+N=7N-$ 1.
A-H Sigmoid
The Sigmoid activation function is defined as:
$\sigma(x)=\frac{1}{1+e^{-x}}$ (A.8)
The total FLOPs for the Sigmoid activation function for a layer with $N$ neuron is $7N$.

### v2

@@ IX-1 Advanced AE architectures
Investigate more advanced AE architectures, such as Bayesian AEs [157], diffusion AEs [158], contractive AEs [159], denoising AEs [160], semi-supervised AEs [161], and masked AEs [162]. Combining and modifying these architectures could lead to improved performance and new applications in communication systems. In addition, it is necessary to make quantitative comparisons of the system performance between different AE architectures. This includes evaluating complexity, efficiency, and effectiveness in various communication scenarios, as well as comparing hardware implementations such as FPGA.
IX-2 Hybrid models
@@
It develops hybrid models that combine the strengths of both data-driven and model-based approaches. This could involve integrating traditional mathematical models with deep learning techniques to enhance the robustness and adaptability of communication systems. The researcher can also investigate the use of quantum circuits in OFDM and MIMO systems. Furthermore, integrating quantum CNNs into existing CNN-based solutions could be a promising direction to handle large block sizes [163].
IX-3 Complexity and Computational Constraints
AE architectures often lead to significant computational and storage demands, which hinders scalability and real-time performance, especially on resource-constrained hardware such as FPGAs [22, 164]. This includes developing low-complexity AE architectures and exploring techniques such as low-precision compression and binary neural networks. Training and inference for large AE models require substantial resources. Even with techniques like tensor decomposition or low-precision methods, complexity can remain a challenge. This requires a carefully adjusted AE structure [112]. Further research can focus on the challenges of real-world deployment, such as the need for extensive training data, the risk of overfitting, and the complexity of practical implementations. This includes developing robust training strategies, data augmentation techniques, and methods for handling noisy or corrupted data.
IX-4 Data and Training Issues
Data Scarcity: collecting sufficient and diverse training data, particularly under varying channel conditions or low SNR, is difficult. Limited data can lead to overfitting and challenges in training models that generalize well. Training Difficulty and Convergence: Training large or complex AE networks can be susceptible to unreliable convergence and entrapment in local minima [29]. Ensuring the learning procedure converges to an optimal solution across diverse channel conditions is challenging [108]. Hyperparameter tuning is often meticulous and critical for performance [14, 53]. Challenges exist related to overfit the training data, which can lead to poor performance with new unseen data. To solve this, a powerful regularization technique is to equip each layer with batch normalization. This normalization significantly accelerates training and effectively mitigates overfitting [165, 50]. Channel Model Dependency and Mismatch: Training models based on inaccurate or incomplete channel representations can lead to suboptimal generalization and performance degradation when deployed on actual channels [24]. Achieving adaptability across varying channel conditions not perfectly matched during training remains a persistent challenge [20].
IX-5 Non-differentiable channels and DSP components
@@
Some approaches require differentiable channel models for back-propagation, which can hinder development unless workarounds are employed [63, 64]. Therefore, it is required to explore alternative approaches for end-to-end learning in real-world scenarios where the channel and transceiver are non-differentiable or the channel is unknown. This could involve reinforcement learning, stochastic perturbation techniques, gradient-free training methods, and genetic algorithms. In addition, differentiable DSP techniques, such as equalization, CPE, and LPN compensation are required. This includes integrating differentiable DSP algorithms into AE models for end-to-end optimization.
IX-6 Hardware and Implementation Issues
It is crucial to incorporate physical constraints, such as channel capacity, power budget, and bandwidth limitations to ensure practical feasibility. However, excessive imposition of constraints can potentially hinder reconstruction performance due to noise in the received data. Therefore, carefully balancing the trade-off between reconstruction loss and physical constraints is vital to achieve optimal system optimization. Hardware effects, such as pulse shaping and quantization, are often not fully accounted for in AE/VAE models, requiring significant effort to implement them as neural network layers [26]. Furthermore, Continuous transmissions pose challenges related to timing synchronization and sampling frequency offset, which must be integrated into the training process. Synchronization in single-carrier AE systems is difficult [26, 27].
IX-7 Scalability and High Dimensions
AE architectures can face the curse of dimensionality, which leads to exponential growth in complexity and training difficulties when dealing with a large number of antennas (MIMO) or high-dimensional data representations. Scaling representations to accommodate large message dimensions can be impractical due to memory and processing demands [13, 64].
IX-8 Interpretability and Trust
A lack of transparency and trust exists in deep learning AE. Understanding why AEs outperform traditional coding technologies remains an open challenge. The complexity of neural networks makes it difficult to provide clear explanations, hindering user trust [166, 31].The interpretability of the AE learning could limit its performance, because existing AE/VAE systems can be perceived as black-box solutions, the complexity of interpreting the performance and behavior of AE/VAE systems is a significant challenges with potential pitfalls in understanding the learned schemes [17, 164, 112].
IX-9 System-Level and Practical Considerations
End-to-end AE designs often lack modularity, which requires separate designs for different source types or scenarios, and there is a lack of coding standards for learned methods. Also there is a need for rate adaptation, as current AEs often have a fixed data rate [17]. In addition, deep Learning AE models must consistently outperform existing solutions in accuracy, efficiency, and robustness for industry integration. The integration of DL into industry standards is not imminent due to these challenges [29].
IX-10 Federated Learning for Distributed AE Training
Researchers can investigate different federated learning algorithms and optimization techniques to enhance the efficiency and convergence speed of distributed AE training [167]. Exploring the use of federated learning in various communication scenarios, such as decentralized communication systems, could provide valuable insights for practical applications. Additionally, addressing challenges such as data heterogeneity, device availability, and communication constraints in federated learning for autoencoders could be important research directions.
IX-11 Edge-AI Integration for Low-Latency Systems
Future research can focus on optimizing the AE models for deployment on edge devices with limited computational resources [168]. Developing lightweight and efficient autoencoder architectures that can run on edge-AI hardware while maintaining high performance could be a key area of interest. Additionally, exploring the integration of edge-AI with other emerging technologies, such as 5G or 6G networks, could further enhance the capabilities of low-latency communication systems.
X Conclusion
@@
The integration of AEs into communication systems represents a paradigm shift from traditional model-based approaches to data-driven methodologies. This survey highlighted the potential of AEs to overcome the inherent limitations of communication systems including mathematical models to provide more accurate and adaptable solutions for next-generation communication systems. The review also highlighted the extensive application of AEs in various domains, including wireless, optical, semantic, and quantum communications, and provided a critical analysis discussion. Despite the progress, challenges persist, such as handling non-differentiable channels, computational complexity, convergence, data representation and data training including data scarcity. Moreover, this paper provides an examination of the computational complexity calculation associated with AE-based systems using FLOPS. This insight is crucial to guide the practical implementation of AEs in resource-constrained environments. Future research should focus on addressing these challenges, exploring advanced AE architectures, and developing hybrid models that combine the strengths of both data-driven and model-based approaches. In doing so, the field can move closer to realizing the full potential of AEs in creating efficient, reliable, and high-performance communication systems.